# HierarchicalPrune: Position-Aware Compression for Large-Scale Diffusion Models

### v2

@@ A.3 Baseline Comparisons
BK-SDM (kim2024bksdm): BK-SDM performs block pruning for U-Net-based diffusion using CLIP (hessel-etal-2021-clipscore) scores as importance metric, where it ranks blocks based on the reduction of CLIP scores of the output images from the model before and after pruning each block. After pruning, BK-SDM distils the compressed model using knowledge from the original unpruned model, to recover performance from the previous steps.
KOALA (lee2024koala): KOALA uses an importance ranking approach by calculating input-output cosine similarity to inform the block pruning strategy for U-Net based DMs. The blocks of the transformer have intrinsic importance relative to input-output transformations by evaluating the cosine similarity between input and output representations/features rather than utilising an external quality metric of transformation.
SANA-Sprint-1.6B (chen2025sanasprintonestepdiffusioncontinuoustime): We use SANA-Sprint-1.6B as the state-of-the-art (SOTA) baseline for an efficient, small-scale diffusion purpose-built model, incorporating techniques such as deep-compression auto-encoder (chen2025deep) and linear attention (linear_attn_katharopoulos2020). Intrinsically, SANA models require training of the transformer from scratch, which departs from compressing existing models.
@@ A.4 Evaluation Framework
Quantitative Metrics: We adopt two representative and complementary image quality assessment metrics, capturing two different aspects of the quality of generation. First, GenEval (ghosh2023geneval) provides an all-encompassing evaluation across three areas, i.e., semantic fidelity, visual fidelity, and prompt adherence. GenEval is now widely used for evaluations of T2I models (xie2025sana; esser2024mmdit). Second, HPSv2 (wu2023human) provides a contrasting perspective on image quality assessment, with demonstrated effectiveness in estimating the perceptual preferences of humans about image quality. HPSv2 has been utilised in several recent works such as KOALA (lee2024koala).
Human Preference Metrics: We conduct a user study to assess human preferences for the generated images that are difficult to capture with the quantitative quality metrics (see Appendix A.5 for further details regarding our user study design). Similarly to (sheynin2023emueditpreciseimage) and (dai2023emuenhancingimagegeneration), we evaluate two aspects: (1) Text Alignment: How well a generated image follows the description of the text prompt, (2) Image Quality: The overall visual quality considering clarity, colour, composition, and other factors.
@@ A.5 Detailed User Study Design
Acknowledging the inconsistency of automated metrics, especially in discerning fine aspects of image quality and preferences for aesthetics, we choose to conduct a full user study of methods following the protocol of (sheynin2023emueditpreciseimage) and (dai2023emuenhancingimagegeneration). In our user evaluation method, we are focused on measuring two aspects of user experience:
@@
Figure 8: Fine-grained contribution analysis of SD3.5 Large Turbo on the HPSv2 dataset by removing either an entire MMDiT block or an intra-block subcomponent. We report the performance drop compared to the original model. (a) An entire MMDiT block is removed following prior depth pruning approaches (lee2024koala; kim2024bksdm; fang2024tinyfusion). (b,c,d,e,f) Each subcomponent of a MMDiT block is removed, revealing the different patterns of importance of each subcomponent.
@@ Appendix B Additional Analysis
Figure 10: Fine-grained contribution analysis of FLUX.1-Schnell on the HPSv2 dataset by removing either an entire MMDiT block or an intra-block subcomponent. We report the performance drop compared to the original model. (a) An entire MMDiT block is removed following prior depth pruning approaches (lee2024koala; kim2024bksdm; fang2024tinyfusion). (b,c,d,e,f,g,h) Each subcomponent of a MMDiT block is removed, revealing the different patterns of importance of each subcomponent. Note that the first part of the model from block 0 to 18 is composed of the transformer blocks with five components (b-f) and the second part of the model is made of single transformer blocks with three components (b,g,h).
@@ C.1 Computational Overhead Analysis
Overall, the contribution analysis phase requires 12.6-27.1 GPU hours for comprehensive architectural profiling. While substantial, this represents a one-time cost that is amortised across all deployments. Furthermore, training small-scale DMs (1-2B parameter-sized) from scratch requires 140,000 and/or 200,000 A100 GPU hours for smaller models such as SD1.4 and SD2.1, respectively, as reported in (2024ICLRgU58d5QeGv_cascade), let alone large-scale DMs (8-11B parameter-sized), which could require significantly more GPU hours to train from scratch. Yet, our contribution analysis cost equals <0.006% of typical DM training.
@@
Overall Computational Requirement: When considered in its entirety, the contribution analysis phase requires 12.6-27.1 GPU hours for comprehensive architectural profiling across different model architectures, and the subsequent distillation pipeline demands 603-1260 A100 GPU hours. While substantial, these represent one-time costs that amortise across all subsequent deployments. To place this computational investment in perspective, training small-scale DMs (1-2B parameters) from scratch requires 140,000-200,000 A100 GPU hours for SD1.4 and SD2.1 models, respectively, as reported (2024ICLRgU58d5QeGv_cascade), let alone large-scale DMs (8-11B parameters), which would require considerably more computational resources. Notably, our contribution analysis cost equals less than 0.006% of typical DM training overhead, while our distillation training process represents less than 0.30-0.63% of the computational cost required to train comparable models from scratch.

### v3

@@ A.3 Baseline Comparisons
BK-SDM (Kim et al. 2024a): BK-SDM performs block pruning for U-Net-based diffusion using CLIP (Hessel et al. 2021) scores as importance metric, where it ranks blocks based on the reduction of CLIP scores of the output images from the model before and after pruning each block. After pruning, BK-SDM distils the compressed model using knowledge from the original unpruned model, to recover performance from the previous steps.
KOALA (Lee et al. 2024): KOALA uses an importance ranking approach by calculating input-output cosine similarity to inform the block pruning strategy for U-Net based DMs. The blocks of the transformer have intrinsic importance relative to input-output transformations by evaluating the cosine similarity between input and output representations/features rather than utilising an external quality metric of transformation.
SANA-Sprint-1.6B (Chen et al. 2025a): We use SANA-Sprint-1.6B as the state-of-the-art (SOTA) baseline for an efficient, small-scale diffusion purpose-built model, incorporating techniques such as deep-compression auto-encoder (Chen et al. 2025b) and linear attention (Katharopoulos et al. 2020). Intrinsically, SANA models require training of the transformer from scratch, which departs from compressing existing models.
@@ A.4 Evaluation Framework
Quantitative Metrics: We adopt two representative and complementary image quality assessment metrics, capturing two different aspects of the quality of generation. First, GenEval (Ghosh et al. 2023) provides an all-encompassing evaluation across three areas, i.e., semantic fidelity, visual fidelity, and prompt adherence. GenEval is now widely used for evaluations of T2I models (Xie et al. 2025; Esser et al. 2024). Second, HPSv2 (Wu et al. 2023a) provides a contrasting perspective on image quality assessment, with demonstrated effectiveness in estimating the perceptual preferences of humans about image quality. HPSv2 has been utilised in several recent works such as KOALA (Lee et al. 2024).
Human Preference Metrics: We conduct a user study to assess human preferences for the generated images that are difficult to capture with the quantitative quality metrics (see Appendix A.5 for further details regarding our user study design). Similarly to (Sheynin et al. 2023) and (Dai et al. 2023), we evaluate two aspects: (1) Text Alignment: How well a generated image follows the description of the text prompt, (2) Image Quality: The overall visual quality considering clarity, colour, composition, and other factors.
@@ A.5 Detailed User Study Design
Acknowledging the inconsistency of automated metrics, especially in discerning fine aspects of image quality and preferences for aesthetics, we choose to conduct a full user study of methods following the protocol of (Sheynin et al. 2023) and (Dai et al. 2023). In our user evaluation method, we are focused on measuring two aspects of user experience:
@@
Figure 8: Fine-grained contribution analysis of SD3.5 Large Turbo on the HPSv2 dataset by removing either an entire MMDiT block or an intra-block subcomponent. We report the performance drop compared to the original model. (a) An entire MMDiT block is removed following prior depth pruning approaches (Lee et al. 2024; Kim et al. 2024a; Fang et al. 2024). (b,c,d,e,f) Each subcomponent of a MMDiT block is removed, revealing the different patterns of importance of each subcomponent.
@@ Appendix B Additional Analysis
Figure 10: Fine-grained contribution analysis of FLUX.1-Schnell on the HPSv2 dataset by removing either an entire MMDiT block or an intra-block subcomponent. We report the performance drop compared to the original model. (a) An entire MMDiT block is removed following prior depth pruning approaches (Lee et al. 2024; Kim et al. 2024a; Fang et al. 2024). (b,c,d,e,f,g,h) Each subcomponent of a MMDiT block is removed, revealing the different patterns of importance of each subcomponent. Note that the first part of the model from block 0 to 18 is composed of the transformer blocks with five components (b-f) and the second part of the model is made of single transformer blocks with three components (b,g,h).
@@ C.1 Computational Overhead Analysis
Overall, the contribution analysis phase requires 12.6-27.1 GPU hours for comprehensive architectural profiling. While substantial, this represents a one-time cost that is amortised across all deployments. Furthermore, training small-scale DMs (1-2B parameter-sized) from scratch requires 140,000 and/or 200,000 A100 GPU hours for smaller models such as SD1.4 and SD2.1, respectively, as reported in (Pernias et al. 2024), let alone large-scale DMs (8-11B parameter-sized), which could require significantly more GPU hours to train from scratch. Yet, our contribution analysis cost equals <0.006% of typical DM training.
@@
Overall Computational Requirement: When considered in its entirety, the contribution analysis phase requires 12.6-27.1 GPU hours for comprehensive architectural profiling across different model architectures, and the subsequent distillation pipeline demands 603-1260 A100 GPU hours. While substantial, these represent one-time costs that amortise across all subsequent deployments. To place this computational investment in perspective, training small-scale DMs (1-2B parameters) from scratch requires 140,000-200,000 A100 GPU hours for SD1.4 and SD2.1 models, respectively, as reported (Pernias et al. 2024), let alone large-scale DMs (8-11B parameters), which would require considerably more computational resources. Notably, our contribution analysis cost equals less than 0.006% of typical DM training overhead, while our distillation training process represents less than 0.30-0.63% of the computational cost required to train comparable models from scratch.